# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v1

@@ 4.6.3 Centralized vs FL :
In the context of comparing the centralized and federated approaches, we meticulously selected parameters conducive to achieving optimal performance for FL and juxtaposed these outcomes with those of the centralized approach. Comparative tests were conducted on diverse datasets, relying on metrics such as accuracy, precision, recall, F1 score, and AUC.
The results were consolidated into stacked bar plots, as depicted in Figure 8, where FL results are represented in orange, and centralized results are represented in blue. It is observed that the outcomes on the NSL-KDD, TON_IOT, UNSW_NB15, and WUSTL-EHMS datasets, presented in Figures 7(a), 7(b), 7(c), and 7(d) respectively, show bars being halved, indicating comparable results between FL and centralized approaches. Consequently, training models on separate data partitions generalizes as effectively as training a global model on the entire dataset. This dual advantage of achieving comparable detection outcomes while preserving data privacy highlights the efficacy and privacy-centric nature of the FL-based IDS, positioning it as a robust and privacy-aware solution for intrusion detection.
@@ 5 DISCUSSION
The proposed framework for intrusion detection in the IoMT system based on DL for anomaly detection, FL for model training and privacy protection, and XAI for enhanced interpretability and explainability. This distinctive approach, in contrast to existing literature, sets our work apart. The introduction of an IDS based on DL underscores its real-time efficiency in swiftly identifying network or host-based attacks that could compromise the integrity, confidentiality, and availability of the IoMT system. By leveraging DL as a ML method, our framework enables the automatic selection of pertinent features, offering an end-to-end solution that operates seamlessly without relying on third-party interventions. This streamlined approach enhances the robustness and autonomy of the intrusion detection process within the IoMT system.
The proposed solution places a paramount emphasis on safeguarding patient privacy by employing FL, which involves sharing model weights instead of actual patient data. By optimizing FL parameters, the communication rounds was significantly reduced, thereby minimizing bandwidth consumption, preventing network congestion, and ensuring the scalability of the system. The distributed nature of FL proves instrumental in positioning the IDS in close proximity to potential attack sources. This proximity enhances the system’s agility, allowing for rapid and effective detection and response to security threats. This decentralized approach not only bolsters the overall security posture but also contributes to the expeditious identification and mitigation of potential risks to patient data within the IoMT ecosystem.
The proposed system delivers an explanation and interpretation of the ML model for anomaly detection, thereby enhancing trust in the capability of the proposed DL-based IDS for anomaly detection. Concurrently, it assists regulators seeking to verify compliance with international standards. For users of the proposed framework, such as patients, trust can be reinforced by presenting a performance history achieved by the system. This historical record serves to strengthen trust in the reliable execution of the system. Moreover, the proposed IDS functions as a valuable decision aid, empowering Chief Information Security Officers (CISOs) to intervene promptly in case of anomaly detection. Whether triggered by communication issues, medical emergencies, or security attacks, this intervention capability ensures a proactive response to safeguard the integrity and security of the IoMT system.
@@
Yet, a significant challenge emerges due to the scarcity of datasets explicitly tailored for IoMT systems, featuring diverse attacks and balanced class instances. This scarcity impedes the validation of security solutions and the comparative analysis of distinct contributions to IoMT system security.
@@ 6 CONCLUSION
In conclusion, this study introduces a framework for an IDS based on an ANN enriched with FL and XAI methods. The synergistic integration of these components enhances robust attack detection in the context of the IoMT, emphasizing data privacy and ensuring model explainability and interpretability. The IDS architecture capitalizes on FL, fostering collaborative model training while upholding the confidentiality of sensitive data, thereby addressing privacy concerns prevalent in healthcare. Additionally, the incorporation of XAI bolsters transparency, ensuring compliance with regulatory requirements and healthcare legislation. This, in turn, cultivates greater trust in the decision-making process of the system among stakeholders. A comprehensive evaluation across diverse datasets containing both network and medical data underscores the applicability and resilience of the proposed solution. Particularly noteworthy are the results showcasing the efficacy of the optimized FL method, achieving an accuracy surpassing 98%, comparable to traditional centralized approaches. Furthermore, the provision of explanations and result interpretations using XAI adds an extra layer of assurance, reinforcing trust for ML model designers, regulators, and users of IoMT systems within the proposed framework.
Future perspectives involve strengthening the solution against poisoning attacks, refining the FL model by improving client selection and dataset partitioning. Additionally, the solution will undergo rigorous testing with newly curated datasets.

### v2

@@ 4.6.3 Centralized vs FL :
In the context of comparing the centralized and federated approaches, parameters are meticulously selected to achieve optimal performance for FL, and these outcomes are compared with those of the centralized approach. Comparative tests are conducted on diverse datasets, relying on metrics such as accuracy, precision, recall, F1 score, and AUC.
The results are consolidated into stacked bar plots, as depicted in Figure 8, where FL results are represented in orange, and centralized results are represented in blue. The outcomes on the NSL-KDD, TON-IOT, UNSW_NB15, and WUSTL-EHMS datasets, presented in Tables 8(a), 8(b), 8(c), and 8(d) respectively, show bars being halved, indicating comparable results between FL and centralized approaches. Consequently, training models on separate data partitions generalizes as effectively as training a global model on the entire dataset. This dual advantage of achieving comparable detection outcomes while preserving data privacy highlights the efficacy and privacy-centric nature of the FL-based IDS, positioning it as a robust and privacy-aware solution for intrusion detection.
5 Comparison with Previous Work
The proposed approach presents several advancements over existing works, including enhanced dataset diversity, FL optimization, XAI integration, and a comprehensive discussion of ethical considerations. Unlike prior studies that rely on a single dataset or focus solely on either network or medical data, this work integrates multiple datasets encompassing both types of data, covering a wide range of attack categories, including threats to confidentiality, integrity, and availability.
Additionally, FL is employed not only for privacy preservation but also with an optimized parameter selection process. This process considers factors such as the number of clients, client fraction, and local epochs, which have not been fully explored in previous FL-based IDS solutions. In terms of explainability, while many existing approaches either lack XAI mechanisms or rely on LIME for local feature explanations, this work incorporates SHAP for global interpretability, complemented by visualization techniques like beeswarm and bar plots, as well as a detection history feature accessible to non-technical users.
Additionally, ethical considerations are explicitly addressed, ensuring compliance with regulatory frameworks like GDPR and HIPAA while incorporating principles of fairness, accountability, and bias mitigation. A comparative summary of these key contributions is provided in Table 6.
Ref Multiple Datasets Diversity of Attack Types FL Implementation FL Parameter Optimization XAI Integration Global Explanations Ethical Discussion
[36] X X X X X X
[37] X X X X X X
[38] X X X
[39] X X X X
[40] X X X X X
[14] X X X X X
[16] X X X X X
[21] X X X X X X X
[24] X X X X X
[29] X X X X X X X
[30] X X X X X X X
[31] X X X X X
[32] X X X X X X
[33] X X X X X
[35] X X X X X X X
Proposed Framework
Table 6: Comparison of the Proposed Approach with Previous Works
@@ 6 DISCUSSION
The proposed framework for intrusion detection in the IoMT system is based on DL for anomaly detection, FL for model training and privacy protection, and XAI for enhanced interpretability and explainability. This distinctive approach, in contrast to existing literature, sets our work apart. The introduction of an IDS based on DL underscores its real-time efficiency in swiftly identifying network or host-based attacks that could compromise the integrity, confidentiality, and availability of the IoMT system. By leveraging DL as a ML method, our framework enables the automatic selection of pertinent features, offering an end-to-end solution that operates seamlessly without relying on third-party interventions. This streamlined approach enhances the robustness and autonomy of the intrusion detection process within the IoMT system.
The proposed solution places a paramount emphasis on safeguarding patient privacy by employing FL, which involves sharing model weights instead of actual patient data. By optimizing FL parameters, the communication rounds was significantly reduced, thereby minimizing bandwidth consumption, preventing network congestion, and ensuring the scalability of the system. The distributed nature of FL proves instrumental in positioning the IDS in close proximity to potential attack sources. This proximity enhances the system’s agility, allowing for rapid and effective detection and response to security threats. This decentralized approach strengthens the overall security posture while contributing to the rapid identification and mitigation of potential risks to patient data within the IoMT ecosystem.
The proposed system provides an explanation and interpretation of the ML model for anomaly detection, thereby enhancing trust in the capability of the proposed DL-based IDS for anomaly detection. Concurrently, it assists regulators seeking to verify compliance with international standards. For users of the proposed framework, such as patients, trust can be reinforced by presenting a performance history achieved by the system. This historical record serves to strengthen trust in the reliable execution of the system. Moreover, the proposed IDS functions as a valuable decision aid, empowering CISO to intervene promptly in case of anomaly detection. Whether triggered by communication issues, medical emergencies, or security attacks, this intervention capability ensures a proactive response to safeguard the integrity and security of the IoMT system.
@@
The proposed framework makes a significant ethical contribution by addressing key concerns related to data privacy, transparency, fairness, and accountability in the context of IoMT systems. By leveraging FL, the framework ensures that sensitive patient data remains on local devices, thereby preserving privacy and complying with stringent regulations such as HIPAA and GDPR. This approach minimizes the risk of data breaches and unauthorized access, fostering trust among patients and healthcare providers. Furthermore, the integration of XAI methods, such as SHAP, enhances the transparency of the decision-making process, allowing stakeholders to understand and verify the model’s predictions. This transparency is crucial for ensuring accountability, as it enables the identification and mitigation of potential biases or errors in the system. Additionally, the framework promotes fairness by using diverse datasets and mitigating biases, ensuring that the benefits of the technology are accessible to all patients without discrimination. These ethical considerations are essential for building trust and ensuring the responsible deployment of AI in healthcare, ultimately contributing to the well-being and safety of patients.
Yet, a significant challenge emerges due to the scarcity of datasets explicitly tailored for IoMT systems, featuring diverse attacks and balanced class instances. This scarcity impedes the validation of security solutions and the comparative analysis of distinct contributions to IoMT system security. Additionally, deploying these solutions in real-world environments introduces unforeseen challenges, such as dynamic network conditions, device heterogeneity, and unpredictable user behavior. Addressing these complexities will be essential to developing robust, scalable, and practical security frameworks for IoMT systems.
@@ 7 CONCLUSION
In conclusion, this study introduces a framework for an IDS based on an ANN enhanced with FL and XAI methods. The synergistic integration of these components enhances robust attack detection in the context of the IoMT, emphasizing data privacy and ensuring model explainability and interpretability. The IDS architecture capitalizes on FL, fostering collaborative model training while upholding the confidentiality of sensitive data, thereby addressing privacy concerns prevalent in healthcare. Additionally, the incorporation of XAI bolsters transparency, ensuring compliance with regulatory requirements and healthcare legislation. This, in turn, cultivates greater trust in the decision-making process of the system among stakeholders. A comprehensive evaluation across diverse datasets containing both network and medical data underscores the applicability and resilience of the proposed solution. Particularly noteworthy are the results showcasing the efficacy of the optimized FL method, achieving an accuracy surpassing 98%, comparable to traditional centralized approaches. Furthermore, the provision of explanations and result interpretations using XAI adds an extra layer of assurance, reinforcing trust for ML model designers, regulators, and users of IoMT systems within the proposed framework. Beyond its technical achievements, the proposed framework makes a significant ethical contribution by ensuring the protection of sensitive patient data through FL, promoting transparency and accountability via XAI, and fostering fairness by mitigating biases in the model. By aligning with international standards such as HIPAA, GDPR, WHO and ISO/IEC 27701, the framework not only enhances the security of IoMT systems but also ensures that the deployment of AI in healthcare is ethical, responsible, and equitable. These ethical considerations are integral to building trust among patients, healthcare providers, and regulators, ultimately contributing to the safe and effective use of AI in healthcare.
Future work will aim to fortify the solution against emerging threats such as poisoning attacks, adversarial ML, and quantum attacks, while optimizing the FL model via advanced client selection and dataset partitioning strategies. Additionally, privacy-preserving approaches, including differential privacy and secure multi-party computation, will be further explored to enhance data confidentiality and security. The solution will undergo extensive real-world testing to identify and address practical challenges, ensuring its robustness and scalability in diverse environments.
8 CRediT authorship contribution statement
Ayoub Si-ahmed: Methodology, Investigation, Writing - Original Draft Mohammed Ali Al-Garadi: Methodology, Writing - Review & Editing, Supervision, Project administration. Narhimene Boustia: Resources, Writing - Review & Editing, Supervision, Project administration.
9 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this paper.